In a nickel bath 30g of Ni are plated after 2 hours with a current intensity of 25 A.
How many liters of H2 will be formed at the same time under standard temperature and pressure?
1
Expert's answer
2019-06-10T04:48:01-0400

When a single nickel atom is deposited, one hydrogen molecule arises; therefore, the number of moles of hydrogen

"\nu=\frac{30}{58.7}=0.51"

then hydrogen volume

"V=\frac{\nu RT}{p}=\frac{0.51\cdot8.31\cdot273}{100} =11.6 (l)."